A 37.80 mH coil carries a constant current of 2.80 A, clockwise. If the current drops to zero and rises in the opposite direction (conterclockwise) to 0.15 A in 93.00 microseconds, what potential difference will appear across the coil during this time?
Answer #1

Potential difference V = LdI/ dt

L = 37.8 mH= 0.0378 H, dI = (2.8) - (-0.15) = 2.95 A , dt = 93ms = 0.093 s

Hence, V = 0.0378 X 2.95 / 0.093 = 1.20 V